Finance. -Problem 11-19

Ellington Electronics wants you to calculate its cost of common stock. During the next 12 months, the company expects to pay dividends (D1) of $2.20 per share, and the current price of its common stock is $40 per share. The expected growth rate is 7 percent. (Do not round intermediate calculations. Round the final answers to 2 decimal places.)

 

a. Compute the cost of retained earnings (Ke).

 

Cost of retained earnings             %

 

b. If a $2.0 flotation cost is involved, compute the cost of new common stock (Kn).

 

Cost of new common stock             %
Answer #1

Expected dividend, \(\mathrm{D} 1=\$ 2.20\) Current price of stock, \(\mathrm{PO}=\$ 40\) Growth rate, \(g=7 \%=0.07\) Cost of retained earnings, \(\mathrm{Kr}=(\mathrm{D} 1 / \mathrm{P} 0)+\mathrm{g}\)

$$ \begin{array}{l} =(\$ 2.20 / \$ 40)+0.07 \\ =0.055+0.07 \\ =0.125 \\ =12.5 \% \end{array} $$

If flotation cost \(=\$ 2\) Per share

Net proceeds from sale of stock, \(\mathrm{Np}=\) Current price - Flotation cost

$$ \begin{array}{l} =\$ 40-\$ 2 \\ =\$ 38 \text { Per share } \end{array} $$

Cost of new common stock, \(\mathrm{Kn}=(\mathrm{D} 1 / \mathrm{Np})+\mathrm{g}\)

$$ \begin{array}{l} =(\$ 2.20 / \$ 38)+0.07 \\ =0.057894737+0.07 \\ =0.127894737 \\ =12.7894737 \% \\ =12.79 \% (Rounded) \end{array} $$

answered by: Conflogy